The maximum range of a bullet fired from a toy pistol mounted on a car at rest is 40m. What will be acute angle of inclination of pistol for max. Range when the car is moving in the direction f firing with uniform velocity v=20m/s, not the horizontal surface?

Grade:12

5 Answers

Akshat Kumar tripathi
6 years ago
We have Range of a projectile = u^2 sin 2theta/g Therefore, u^2 sin 2 theta /g = 40400*sin 2 theta /10 = 40 Hence, theta = 45 degree
K
5 years ago
     u^2/g = 100
=> u = 20 m/s
 
t = 2ugsina and R = (ucosa+20)(2usina/g)
Put dR/da = 0 => cosa = ½
=> a = 60°.
Deepika Pant
5 years ago
The initial velocity,
Anagha
4 years ago
when at rest, the range of the projectile is usin(2x) / g. This is maximum when sin(2x)=1.
So, 
40 = u/ g => u = 20
Now, as the car is moving with 20 m/s, the initial velocity of the projectile in the x direction is (20 + u cos(x)). 
As acceleration along the x direction is zero, range is T * V
=> (2u sin(x) / g)* (20 + u cos(x))
on differentiating and equating to zero (to get the maximum) we get x =60 
 
Hope that helps! 
 
ankit singh
3 years ago

gu2=100
u=20 m/s
t=g2usinθ
and R=(ucosθ+20)(g2usinθ)
Put dθdR=0
cosθ=1/2
θ=60.

ankit singh nt bhpal
thanks and regards